# Harnessing Batched BLAS/LAPACK Kernels on GPUs for Parallel Solutions of Block Tridiagonal Systems

### v1

@@ Abstract
We present a GPU implementation for the factorization and solution of block-tridiagonal symmetric positive definite (SPD) linear systems, which commonly arise in time-dependent estimation and optimal control problems. Our method employs a recursive algorithm based on Schur complement reduction, enabling the transformation of the system into a hierarchy of smaller, independent blocks that can be efficiently solved in parallel using batched BLAS/LAPACK routines. While batched routines have been utilized in sparse solvers [22], our approach applies these kernels in a tailored way by exploiting the block-tridiagonal structure known a priori. Performance benchmarks based on our open-source, cross-platform (NVIDIA and AMD) implementation, TBD-GPU, demonstrate the advantages of tailored utilization of batched routines: achieving substantial speed-ups compared to state-of-the-art CPU direct solvers, including CHOLMOD and HSL MA57, while remaining competitive with NVIDIA cuDSS. However, our current implementation still performs sequential calls of batched routines at each recursion level, and the block size must be sufficiently large to adequately amortize the kernel launch overhead.
@@ 1 Introduction
  • Kalman Smoothing: Kalman smoothing is a classical technique used to estimate the state of a dynamical system based on noisy measurements over a given time horizon. Under the linear-Gaussian setting, the maximum a posteriori (MAP) estimation problem is formulated such that the prediction error from the system dynamics and the measurement error are penalized quadratically within the objective function. The Rauch–Tung–Striebel (RTS) smoother [21], in particular, yields a block-tridiagonal SPD Hessian matrix [4], enabling the estimation problem to be efficiently solved using specialized block-tridiagonal solvers. Batched GPU solvers hold significant potential for facilitating large-scale state estimation in applications such as high-energy physics track fitting [3].
  • Model predictive control (MPC): The MPC methods make control decisions by solving optimal control problems in real-time. Due to the presence of inequality constraints, these problems are often addressed using interior-point method (IPM), where most of the computation time is dedicated to solving the associated KKT systems [14]. These KKT systems become increasingly ill-conditioned as the iterate approaches the solution because of the active inequality constraints; therefore, direct linear solvers are typically employed due to their numerical robustness. The KKT systems for the control problems exhibit a block tridiagonal structure resulting from the temporal coupling of system dynamics over the prediction horizon. Although the original KKT systems are often indefinite, they can be transformed into SPD systems through the application of condensation strategies [19]. GPU acceleration presents a promising avenue for achieving the performance required for real-time control [19, 20].
@@
To address this limitation, this paper presents a recursive Schur complement-based strategy, a technique also known as substructuring in the domain decomposition literature [24, 25] (described in Figure 1.1). Similar formulations have been explored in the context of block-tridiagonal systems [6], but without exploiting batched BLAS/LAPACK kernels, which are the central focus of our work. In particular, a symmetric permutation is employed to reorganize the matrix into a substantial set of independent interior blocks, along with a smaller set of coupling separator variables. This reordering facilitates the computationally intensive elimination of all interior segments, allowing it to be executed in a single, massively parallel stage using batched kernels. The resulting smaller, denser, and still block-tridiagonal Schur complement system focuses solely on the separators, which are then solved by recursively applying the same procedure. This recursive approach systematically transforms a sparse, latency-bound problem into a hierarchy of dense, compute-bound kernels that are well-suited for GPU execution. Furthermore, given that the block-tridiagonal structure is known a priori, we can perform tailored reordering that leverages this structure, thereby enhancing GPU utilization. We will later present a tailored recursive partitioning strategy that can maximize the utilization of parallel computing capabilities and achieve favorable scaling results for parallel execution.
The proposed strategy specifically aims to leverage batched basic linear algebra subprograms (BLAS)/linear algebra package (LAPACK) in an optimized manner. Batched variants of BLAS and LAPACK routines hold significant promise for accelerating the solution of these systems, as demonstrated in [22]. These operations perform identical linear algebra operations across a large number of independent small to medium-sized matrices, effectively eliminating kernel launch overhead and maximizing GPU utilization. The recursive Schur-complement-based strategy described above can be implemented using three standard batched BLAS/LAPACK routines: Cholesky factorization (potrf), triangular solve (trsm), and matrix-matrix multiplication (gemm), all of which are widely available in vendor-specific or open-source GPU libraries. All these kernels rely on multiply-accumulate operations at their core, making them particularly well-suited to exploit the high computational throughput of modern GPUs, including tensor cores. Furthermore, since we target SPD systems, numerical pivoting techniques typically used in getrf or sytrf kernels—which can potentially slow down computation—are unnecessary. Thus, the factorization phase remains highly parallelizable and efficient.
@@
The recursive Schur complement strategy described above is implemented in our open-source solver, TBD-GPU. The solver takes in the block-tridiagonal components as inputs and offers numerical factorization and solve capabilities (with support for multiple right-hand sides). Leveraging the multiple dispatch feature [7], our solver is designed to seamlessly switch between CUDA and ROCm backends while also supporting both single and double precision floating-point arithmetic.
@@ 1.1 Related Work
The Thomas algorithm is a serial algorithm for solving tridiagonal systems [26] (i.e., $n=1$ in Eq. 1.1) with $O(N)$ time complexity. To overcome this bottleneck, parallel algorithms such as cyclic reduction (CR) [11] and block cyclic reduction [10] were developed. While these methods expose significant parallelism, it often comes at the cost of increased arithmetic operations or reduced numerical stability. A comprehensive survey of these methods is provided by Gander and Golub [9]. Block-tridiagonal extensions of these algorithms are critical in domains like Kalman smoothing and optimal control, where block sizes may be small but the problem horizon $N$ is large. Although stable block elimination approaches are well established [5], developing efficient and robust formulations for massively parallel architectures remains an active area of research. Note that block variants of the Thomas algorithms and cyclic reduction methods are special cases of the recursive Schur complement framework presented in this paper. Our Schur complement method allows for a trade-off between arithmetic complexity and enhanced parallelism and GPU utilization.
@@
The implementation of GPU solvers can largely benefit from vendor-provided dense linear algebra libraries. NVIDIA provides cuBLAS for dense operations, cuSOLVER for factorizations, and cuSPARSE/cuDSS for sparse direct methods [15, 17, 18, 16]. AMD offers similar functionality through its ROCm ecosystem [2]. The routines available in these libraries are highly optimized at the hardware architecture level, and thus, enables high-performance linear algebra operations. The key enabler for solving many small to medium-sized structured problems is the emergence of batched BLAS and LAPACK kernels. These routines amortize kernel launch overhead and maintain high GPU utilization by executing the same operation across thousands of independent matrices. This is ideal for problems that are too small to individually saturate the hardware. The Batched BLAS (BBLAS) interface proposal [1], along with implementations in MAGMA [28] and vendor libraries, has made these powerful primitives widely accessible. Our algorithm leverages these batched kernels to maximize high-throughput operations during block eliminations.
@@
Several libraries specifically target tridiagonal or block-tridiagonal systems on GPUs. TridiGPU offers bandwidth-optimized implementations of the Thomas algorithm and PCR variants [13], while PaScaL_TDMA implements a hybrid Thomas/PCR approach for CUDA [12]. The PfSolve framework generalizes cyclic reduction strategies across multiple architectures, including GPUs [27]. AMD’s ROCm library contains a block-tridiagonal factorization routines [2], but as of this writing (ROCm v6.4.1) , this routine currently do not work for all problem sizes. These solvers achieve excellent performance for scalar tridiagonal or narrow-banded systems, but they generally do not leverage the modern GPU architecture like tensorcores and support the large block sizes or hierarchical Schur complement factorizations required for handling a large number of blocks. Our work addresses this gap by combining stable block elimination with recursive batching and GPU-optimized kernels.
@@ 1.2 Contributions
First, we introduce a recursive Schur-complement strategy optimized for block-tridiagonal systems, focusing on leveraging batched BLAS/LAPACK operations on GPUs. While nested Schur complement strategies have been explored in the literature [29], their application has been confined mainly to dynamic programming contexts and has yet to be studied in relation to SPD block-tridiagonal systems. Furthermore, the algorithm has not been contextualized within the framework of batched BLAS/LAPACK operations. Although the use of batched BLAS/LAPACK operations in general sparse direct solvers has been investigated previously [22], our work specifically tailors the algorithm to the block-tridiagonal structure, enabling the implementation of optimized reordering strategies that enhance maximal GPU utilization.
@@ 4.1 Experimental Setup
We benchmark across fixed total problem size $N\times n=262{,}144$, sweeping $n\in\{32,64,128,256,512,1024\}$ (and corresponding $N$). Table 4.1 summarizes performance in FP64. Our code is available in the link specified in the previous footnote.
@@ 4.2 Discussion of Results
Figure 4.1: Kalman smoothing with rotational latent dynamics (n=256n{=}256, m=1024m{=}1024, N=100N{=}100). We plot the ground truth (solid) and smoothed trajectory (dash) in a representative 2D oscillatory subspace.
@@ 4.3 Application: Kalman Smoothing
where the matrices $H$, $R$, $Q$, $G$, and the vector $\zeta$ are defined as in [4]. The corresponding optimality conditions yield an SPD block-tridiagonal system:
@@
To showcase the solver in a realistic high-dimensional regime, we draw inspiration from neural population dynamics in motor cortex, where latent activity exhibits smooth rotational structure and is observed through a large number of noisy spike trains [23, 8]. We instantiate the model with latent state dimension $n=256$, observation dimension $m=1024$, and horizon $N=100$ (10 s sampled at $\Delta t=0.1$ s).
@@
On this tall-observation, large-state problem, TBD-GPU avoids general sparse overhead and achieves a clear runtime advantage over cuDSS. Figure 4.1 illustrates the recovered latent trajectory: the smoothed estimates (dashed) closely track the true latent dynamics (solid) across two representative coordinates in the rotational subspace, confirming phase and amplitude alignment.

### v2

@@ Abstract
We present a GPU implementation for the factorization and solution of block-tridiagonal symmetric positive definite (SPD) linear systems, which commonly arise in time-dependent estimation and optimal control problems. Our method employs a recursive algorithm based on Schur complement reduction, enabling the transformation of the system into a hierarchy of smaller, independent blocks that can be efficiently solved in parallel using batched BLAS/LAPACK routines. While batched routines have been utilized in sparse solvers [rennichAcceleratingSparseCholesky2014], our approach applies these kernels in a tailored way by exploiting the block-tridiagonal structure known a priori. Performance benchmarks based on our open-source, cross-platform (NVIDIA and AMD) implementation, TBD-GPU, demonstrate the advantages of tailored utilization of batched routines: achieving substantial speed-ups compared to state-of-the-art CPU direct solvers, including CHOLMOD and HSL MA57, while remaining competitive with NVIDIA cuDSS. However, our current implementation still performs sequential calls of batched routines at each recursion level, and the block size must be sufficiently large to adequately amortize the kernel launch overhead.
@@ 1 Introduction
  • Kalman Smoothing: Kalman smoothing is a classical technique used to estimate the state of a dynamical system based on noisy measurements over a given time horizon. Under the linear-Gaussian setting, the maximum a posteriori (MAP) estimation problem is formulated such that the prediction error from the system dynamics and the measurement error are penalized quadratically within the objective function. The Rauch–Tung–Striebel (RTS) smoother [Rauch1965], in particular, yields a block-tridiagonal SPD Hessian matrix [ARAVKIN2021821], enabling the estimation problem to be efficiently solved using specialized block-tridiagonal solvers. Batched GPU solvers hold significant potential for facilitating large-scale state estimation in applications such as high-energy physics track fitting [Ai2021].
  • Model predictive control (MPC): The MPC methods make control decisions by solving optimal control problems in real-time. Due to the presence of inequality constraints, these problems are often addressed using interior-point method (IPM), where most of the computation time is dedicated to solving the associated KKT systems [nocedalNumericalOptimization2006]. These KKT systems become increasingly ill-conditioned as the iterate approaches the solution because of the active inequality constraints; therefore, direct linear solvers are typically employed due to their numerical robustness. The KKT systems for the control problems exhibit a block tridiagonal structure resulting from the temporal coupling of system dynamics over the prediction horizon. Although the original KKT systems are often indefinite, they can be transformed into SPD systems through the application of condensation strategies [pacaudGPUacceleratedDynamicNonlinear2024]. GPU acceleration presents a promising avenue for achieving the performance required for real-time control [pacaudGPUacceleratedDynamicNonlinear2024, du2025gatogpuacceleratedbatchedtrajectory].
@@
To address this limitation, this paper presents a recursive Schur complement-based strategy, a technique also known as substructuring in the domain decomposition literature [Smith1996, TEREKHOV2013245] (described in Figure 1.1). Similar formulations have been explored in the context of block-tridiagonal systems [belov2017], but without exploiting batched BLAS/LAPACK kernels, which are the central focus of our work. In particular, a symmetric permutation is employed to reorganize the matrix into a substantial set of independent interior blocks, along with a smaller set of coupling separator variables. This reordering facilitates the computationally intensive elimination of all interior segments, allowing it to be executed in a single, massively parallel stage using batched kernels. The resulting smaller, denser, and still block-tridiagonal Schur complement system focuses solely on the separators, which are then solved by recursively applying the same procedure. This recursive approach systematically transforms a sparse, latency-bound problem into a hierarchy of dense, compute-bound kernels that are well-suited for GPU execution. Furthermore, given that the block-tridiagonal structure is known a priori, we can perform tailored reordering that leverages this structure, thereby enhancing GPU utilization. We will later present a tailored recursive partitioning strategy that can maximize the utilization of parallel computing capabilities and achieve favorable scaling results for parallel execution.
The proposed strategy specifically aims to leverage batched basic linear algebra subprograms (BLAS)/linear algebra package (LAPACK) in an optimized manner. Batched variants of BLAS and LAPACK routines hold significant promise for accelerating the solution of these systems, as demonstrated in [rennichAcceleratingSparseCholesky2014]. These operations perform identical linear algebra operations across a large number of independent small to medium-sized matrices, effectively eliminating kernel launch overhead and maximizing GPU utilization. The recursive Schur-complement-based strategy described above can be implemented using three standard batched BLAS/LAPACK routines: Cholesky factorization (potrf), triangular solve (trsm), and matrix-matrix multiplication (gemm), all of which are widely available in vendor-specific or open-source GPU libraries. All these kernels rely on multiply-accumulate operations at their core, making them particularly well-suited to exploit the high computational throughput of modern GPUs, including tensor cores. Furthermore, since we target SPD systems, numerical pivoting techniques typically used in getrf or sytrf kernels—which can potentially slow down computation—are unnecessary. Thus, the factorization phase remains highly parallelizable and efficient.
@@
The recursive Schur complement strategy described above is implemented in our open-source solver, TBD-GPU. The solver takes in the block-tridiagonal components as inputs and offers numerical factorization and solve capabilities (with support for multiple right-hand sides). Leveraging the multiple dispatch feature [Bezanson2017Julia], our solver is designed to seamlessly switch between CUDA and ROCm backends while also supporting both single and double precision floating-point arithmetic.
@@ 1.1 Related Work
The Thomas algorithm is a serial algorithm for solving tridiagonal systems [Thomas1949] (i.e., $n=1$ in Eq. 1.1) with $O(N)$ time complexity. To overcome this bottleneck, parallel algorithms such as cyclic reduction (CR) [hockney1965cr] and block cyclic reduction [sweet1977] were developed. While these methods expose significant parallelism, it often comes at the cost of increased arithmetic operations or reduced numerical stability. A comprehensive survey of these methods is provided by Gander and Golub [GanderGolub1997CyclicReduction]. Block-tridiagonal extensions of these algorithms are critical in domains like Kalman smoothing and optimal control, where block sizes may be small but the problem horizon $N$ is large. Although stable block elimination approaches are well established [Aravkin2014], developing efficient and robust formulations for massively parallel architectures remains an active area of research. Note that block variants of the Thomas algorithms and cyclic reduction methods are special cases of the recursive Schur complement framework presented in this paper. Our Schur complement method allows for a trade-off between arithmetic complexity and enhanced parallelism and GPU utilization.
@@
The implementation of GPU solvers can largely benefit from vendor-provided dense linear algebra libraries. NVIDIA provides cuBLAS for dense operations, cuSOLVER for factorizations, and cuSPARSE/cuDSS for sparse direct methods [NVIDIACUBLAS2025, NVIDIACUSOLVER2025, NVIDIACUSPARSE2025, NVIDIACUDSS2025]. AMD offers similar functionality through its ROCm ecosystem [ROCm2025]. The routines available in these libraries are highly optimized at the hardware architecture level, and thus, enables high-performance linear algebra operations. The key enabler for solving many small to medium-sized structured problems is the emergence of batched BLAS and LAPACK kernels. These routines amortize kernel launch overhead and maintain high GPU utilization by executing the same operation across thousands of independent matrices. This is ideal for problems that are too small to individually saturate the hardware. The Batched BLAS (BBLAS) interface proposal [dongarra2017bblas], along with implementations in MAGMA [MAGMA2010] and vendor libraries, has made these powerful primitives widely accessible. Our algorithm leverages these batched kernels to maximize high-throughput operations during block eliminations.
@@
Several libraries specifically target tridiagonal or block-tridiagonal systems on GPUs. TridiGPU offers bandwidth-optimized implementations of the Thomas algorithm and PCR variants [klein2023tridigpu], while PaScaL_TDMA implements a hybrid Thomas/PCR approach for CUDA [kkpc2020]. The PfSolve framework generalizes cyclic reduction strategies across multiple architectures, including GPUs [tolmachev2025pfsolve]. AMD’s ROCm library contains a block-tridiagonal factorization routines [ROCm2025], but as of this writing (ROCm v6.4.1) , this routine currently do not work for all problem sizes. These solvers achieve excellent performance for scalar tridiagonal or narrow-banded systems, but they generally do not leverage the modern GPU architecture like tensorcores and support the large block sizes or hierarchical Schur complement factorizations required for handling a large number of blocks. Our work addresses this gap by combining stable block elimination with recursive batching and GPU-optimized kernels.
@@ 1.2 Contributions
First, we introduce a recursive Schur-complement strategy optimized for block-tridiagonal systems, focusing on leveraging batched BLAS/LAPACK operations on GPUs. While nested Schur complement strategies have been explored in the literature [wrightPartitionedDynamicProgramming1991], their application has been confined mainly to dynamic programming contexts and has yet to be studied in relation to SPD block-tridiagonal systems. Furthermore, the algorithm has not been contextualized within the framework of batched BLAS/LAPACK operations. Although the use of batched BLAS/LAPACK operations in general sparse direct solvers has been investigated previously [rennichAcceleratingSparseCholesky2014], our work specifically tailors the algorithm to the block-tridiagonal structure, enabling the implementation of optimized reordering strategies that enhance maximal GPU utilization.
@@ 4.1 Experimental Setup
Figure 4.1: Scaling performance of GPU batched solver. Top: factorize and solve times vs. block size $n$ (fixed $N=64$). Bottom: factorize and solve times vs. number of blocks NN (fixed $n=64$).
@@
We benchmark across fixed total problem size $N\times n=262{,}144$, sweeping $n\in\{32,64,128,256,512,1024\}$ (and corresponding $N$). Table 4.1 summarizes performance in FP64. Figure 4.1 shows scaling results for the GPU batched solver. The top panel varies the block size $n$ at fixed number of blocks $N=64$, while the bottom panel varies the number of blocks $N$ at fixed block size $n=64$. Our code is available in the link specified in the previous footnote.
@@ 4.2 Discussion of Results
Scaling behavior.
In Figure 4.1, with $N$ fixed, the total work per step follows dense block costs (factorization $\Theta(n^{3})$ and solves $\Theta(n^{2})$) and the wall-time curves grow accordingly, with factorization dominant. With $n$ fixed, the total work increases proportionally to $N$, but the wall time grows sublinearly due to the multilevel recursive algorithm.
@@ 4.3 Application: Kalman Smoothing
where the matrices $H$, $R$, $Q$, $G$, and the vector $\zeta$ are defined as in [ARAVKIN2021821]. The corresponding optimality conditions yield an SPD block-tridiagonal system:
@@
To showcase the solver in a realistic high-dimensional regime, we draw inspiration from neural population dynamics in motor cortex, where latent activity exhibits smooth rotational structure and is observed through a large number of noisy spike trains [sabatini2024reach, churchland2012neural]. We instantiate the model with latent state dimension $n=256$, observation dimension $m=1024$, and horizon $N=100$ (10 s sampled at $\Delta t=0.1$ s).
@@
On this tall-observation, large-state problem, TBD-GPU avoids general sparse overhead and achieves a clear runtime advantage over cuDSS. We include figures about the recovered latent trajectory in Appendix A.
@@ 5 Conclusions
Acknowledgement
We gratefully acknowledge generous funding provided by Charles Cahn. We also thank Ryan Senne for valuable discussions on the neuroscience applications of this work.
A Kalman Smoothing Plots
Figure A.1: Kalman smoothing with rotational latent dynamics (n=256n{=}256, m=1024m{=}1024, N=100N{=}100). We plot the ground truth (solid) and smoothed trajectory (dash) in a representative 2D oscillatory subspace.
Discussion.
This example highlights the qualitative behavior of our smoother on a high-dimensional latent system. Figure A.1 illustrates the recovered latent trajectory: the smoothed estimates (dashed) closely track the true latent dynamics (solid) across two representative coordinates in the rotational subspace, confirming phase and amplitude alignment.